OpenStudy (anonymous):

Does anybody know a relatively intuitive method to integrate 1/cos(x) ? Thanks!

OpenStudy (diyadiya):

\[\int\limits secxdx= \log|secx+tanx|+C\]

OpenStudy (diyadiya):

@Mimi_x3 You were right =)

OpenStudy (mimi_x3):

Write it as: \[\int\limits secxdx \] Then.. \[\int\limits\frac{secx*secx+tanx}{1*secx+tanx} \] I was wrong diya, lol.

OpenStudy (mimi_x3):

lol, sorry. \[\large \int\limits\frac{\sec^{2}x+secxtanx}{secx+tanx} \] happy now? :P

OpenStudy (anonymous):

Thanks, all of you :) Mimi's way is quite nice, but is there no way that doesn't involve memorising a "trick" like that one?

OpenStudy (diyadiya):

\[\int\limits\limits\frac{sec^2x+secxtanx}{secx+tanx}\] Let Secx+tanx =u \[du/dx=secxtanx+\sec^2x\] \[\int\limits \frac{du}{u}=\log|u|+C=\log|secx+tanx|+C\]

OpenStudy (anonymous):

See http://www.math.com/tables/integrals/more/csc.htm

OpenStudy (amistre64):

the integration for sec(x) was discovered prior to the calculus being formed.

OpenStudy (amistre64):

Mercator and mapmaking are involved.